OpenStudy (margotlakus):

Can someone help me by explaining how to solve logarithmic differentiation?

OpenStudy (mathmale):

Happy to try, but would MUCH prefer to work on an actual example. Otherwise your question is much too broad. By the way, one does not "solve" log differentiation; one "applies" it.

OpenStudy (margotlakus):

thank you! \[-3\ln (8x)\]

OpenStudy (mathmale):

I'm assuming that you have the function f(x)=-3ln (8x). One approach (not the only approach) would be to re-write this function as f(x)=-3 [ln 8 + ln x]. ln 8 is a constant. What is its derivative? x is a variable. What is the derivative of ln x? What is the derivative of ln 8x? Multiply this result by -3 and you'll be done. You must label your result: dy/dx = -3 [ ? ]

OpenStudy (margotlakus):

That's what I don't understand, the derivative of ln8x & lnx

OpenStudy (mathmale):

Another way would be to use the chain rule. 8x is a function in its own right. Thus, the derivative with respect to x of f(x) = -3 ln 8x would be\[f'(x)=-3\frac{ d }{ dx }\ln (8x)\]

OpenStudy (mathmale):

Note that if u is a separate function, then the derivative of y=ln u is\[\frac{ dy }{ dx }=\frac{ 1 }{ u }\frac{ du }{ dx }\]

OpenStudy (mathmale):

Try it.

OpenStudy (margotlakus):

So would the final answer be \[\frac{ dy }{ dx }=-3[\frac{ 1 }{ 8x }]\] ?

zepdrix (zepdrix):

Logarithmic differentiation is usually applied to a problem involving an exponential. Was this your original function? \(\large\rm y=(8x)^{-3}\) and you took the log of each side? \(\large\rm ln(y)=-3ln(8x)\) Or you're just simply trying to find the derivative of the log function?

OpenStudy (margotlakus):

I was given \[f(x)=-3\ln (8x)\]

zepdrix (zepdrix):

Ah ok fair enough :)\[\large\rm \frac{d}{dx}-3\ln(8x)=-3\frac{1}{8x}\frac{d}{dx}(8x)\]Don't forget your all-important `chain rule`! Multiply by the derivative of the inner function.

OpenStudy (margotlakus):

Oh my goodness thank you!